Solve using systems of equations:
x+9=(460/4)
-(460/4) is a fraction.
-When i tried this intitally, I got x=115, but it was marked wrong.

Found 2 solutions by amarjeeth123, DrBeeee:
Answer by amarjeeth123(570)

x+9=(460/4)
x+9=115
x=106
The value of x is 106.

Answer by DrBeeee(684)

You forgot to subtract 9 from each side to get x = 106.
Given:
(1) x + 9 = 460/4 or
(2) x + 9 = 115
Now subtract 9 from each side to get
(3) x + 9 - 9 = 115 - 9 or
(4) x + 0 = 106 or
(5) x = 106